S7.17 Markland Manufacturing intends to increase capacity by overcoming a bottleneck operation by adding new equipment. Two vendors have presented proposals. The fixed costs for proposal A are $50 000, and for proposal B, $70 000. The variable cost for A is $12.00, and for B, $10.00. The revenue generated by each unit is $20.00. a) What is the break-even point in units for proposal A?
b) What is the break-even point in units for proposal B? P
Answer #1

a) let the break even point be X, then

Total cost = Total Revenue

50000 + 12*X = 20*X

8*X = 50000

X = 6250 Units

b) Let the break even point be Y, then

70000 + 10*Y = 20*Y

10*Y = 70000

Y = 7000 Units